# Machine Learning Time Propagators for Time-Dependent Density Functional Theory Simulations

### v2

@@ 1 Introduction
Time-Dependent Density Functional Theory (TDDFT) [1] is a widely used method to study the evolution of electronic structure under the influence of external time-dependent potentials. It is used to calculate various excited state properties such as excitation energies [2], charge transfer [3], stopping power [4], optical absorption spectra [5] and non-linear optical properties [6]. Due to its favorable balance between accuracy and computational cost, TDDFT has been applied in various domains including photocatalysis [7], biochemistry [8], nanoscale devices [9] and the study of light-matter interactions in general [10].
@@
For weak perturbations, the linear response formalism of TDDFT is used to calculate the excitation spectrum of a system  [11]. It is calculated using the Casida equation [12] as an eigenvalue problem that describes the first order response of the density. In contrast, the electron density is directly propagated in time under the real time formalism. Real time TDDFT can be used to calculate the nonlinear response of the density under strong perturbations such as ultrafast electron dynamics with strong laser fields [13].
@@
There are multiple components that must be decided for real time TDDFT calculations. These include preparation of the initial density and orbitals, choice of the exchange-correlation functional, the form of the external time-dependent potential which is determined by the problem to be solved and the choice of the time-propagation scheme. The time evolution is a significant fraction of the computational cost [14, 15].
@@
Machine learning (ML) for accelerating scientific simulations is a rapidly growing area of research [16]. A variety of unsupervised and data-driven models have been developed to solve differential equations across a wide range of domains [17].
@@
Neural operators (NOs) [18] are a class of models that map function-to-function spaces, as opposed to the finite-dimensional vector space mappings of neural networks. This is especially useful for partial differential equation (PDE) problems where experimental or simulation data is available. Fourier neural operators (FNOs) [19] are a type of NO which represent operator weights in Fourier space. The main advantages of FNOs are that they generalize well across function spaces, and being resolution-invariant, they can be used for inference on higher resolution grids than the training set grids. FNOs have also been used for forward and inverse PDE problems in various domains [20].
@@
While many ML models have been developed for the electronic structure in ground state DFT [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], there have been relatively fewer efforts focused on TDDFT. Some applications of ML for TDDFT include development of exchange-correlation potentials [37, 38, 39, 40] and predicting properties such as spectra [41, 42] and stopping power [43]. In addition, there is prior work on developing ML models for quantum dissipative systems [44, 45, 46] and for FNO based TDDFT time propagators for model systems under the influence of a fixed laser field [47].
@@ 2.1 Time-Dependent Density Functional Theory
where $\underline{\bm{r}}$ represents the collective coordinates of the $N$ electrons, $\Psi(\underline{\bm{r}},t)$ is the many-body wavefunction and $\hat{H}$ is the Hamiltonian operator, representing the total energy of the system. For this problem the form of $\hat{H}$ is $\hat{H}(t)=\hat{T}+\hat{W}+\hat{V}(t)$, where $\hat{T}=\sum_{j=1}^{N}-\nabla_{j}^{2}/2$ is the kinetic energy operator, $\hat{W}=\sum_{1\leq j<k\leq N}\frac{1}{\left|\mathbf{r}_{j}-\mathbf{r}_{k}\right|}$ represents the electron-electron interaction, and $\hat{V}(t)=v_{\mathrm{{\scriptscriptstyle ext}}}(\bm{r},t)$ is the time-dependent external potential operator which includes the static external potential due to the ions and an external time-dependent potential that drives the system. Note that in writing this Schrödinger equation we are assuming the Born-Oppenheimer approximation, i.e. we are treating the dynamics of the ions as occurring on a much longer time scale than the dynamics of the electrons, and can therefore treat the ions as classical point-like particles. The initial state $\Psi(\underline{\bm{r}})=\Psi(\underline{\bm{r}},t_{0})$ is obtained by solving the time-independent Schrödinger equation
@@
which is an eigenvalue problem where $E$ denotes the eigenvalues and corresponds to the total energy of the system.
@@
TDDFT is based on the Runge-Gross theorem [1] which states that for a system with a given ground state many-body wavefunction $\Psi_{0}=\Psi_{0}\left(\boldsymbol{r},t_{0}\right)$, there exists a unique mapping between the potential and the time-dependent density. The density of the interacting many-body system can be obtained by solving a system of fictitious non-interacting particles governed by the time-dependent Kohn-Sham (TDKS) equations [48]
@@
$v_{\mathrm{{\scriptscriptstyle S}}}[n](\bm{r},t)=v_{\mathrm{{\scriptscriptstyle ext}}}(\bm{r},t)+v_{\mathrm{{\scriptscriptstyle H}}}[n](\bm{r},t)+v_{\mathrm{{\scriptscriptstyle XC}}}[n](\bm{r},t),$ (6)
@@
While in theory, $v_{\mathrm{{\scriptscriptstyle XC}}}(\bm{r},t)$ depends on the density at all previous time steps, the adiabatic approximation is often used in practice [49]. We use the adiabatic local density approximation (ALDA). Under ALDA, $v_{\mathrm{{\scriptscriptstyle XC}}}(\bm{r},t)$ is approximated as:
$v_{\mathrm{{\scriptscriptstyle XC}}}^{\text{ALDA}}[n,\Psi_{0},\Phi_{0}](\mathbf{r},t)=v_{\mathrm{{\scriptscriptstyle XC}}}^{\text{LDA}}[n](\mathbf{r},t),$ (7)
@@
These coupled equations are solved through iterative numerical algorithms [14].
@@ 2.1.1 Time Propagators
Numerous methods have been developed and analyzed for approximating the short-time propagator $\hat{U}(t_{j}+\Delta t,t_{j})$ [14, 15], including Magnus expansions that approximate the exponent in $\hat{U}(\Delta t)=\exp(\hat{\Omega}(\Delta t))$, split-operator schemes effective when the Hamiltonian separates into analytically solvable parts (e.g., kinetic and potential), explicit and implicit Runge–Kutta methods adapted to the Schrödinger equation, exponential-midpoint and other exponential integrators, and implicit schemes such as Crank–Nicolson [50]. The choice of propagator often involves a trade-off between accuracy, stability (especially concerning the maximum usable time step $\Delta t$), computational cost per step, and conservation properties.
For this work, the reference data for training our neural network model was generated using the Crank-Nicolson method [50]. This is an implicit method, formulated as:
@@ 2.1.2 Model Systems
We use the LDA functional due to its simplicity and accuracy for such simple one-dimensional systems. The density evolution is calculated by solving the time-dependent Kohn-Sham equations with the external potential $v_{\text{ext}}(x,t)$ using the Crank-Nicholson scheme. The numerical simulation is performed using the Octopus real time TDDFT code [51].
@@ 2.2 Physics-Informed Machine Learning
Neural operators (NOs) [52, 53] extend neural networks by mapping functions to functions instead of finite-dimensional vectors. While a neural network maps $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$, a neural operator maps function spaces, $\mathcal{G}:\mathcal{A}\rightarrow\mathcal{U}$. Our goal is to approximate the non-linear map $\mathcal{G}^{\dagger}:\mathcal{A}\rightarrow\mathcal{U}$ with a neural operator $\mathcal{G}_{\theta}$, parameterized by $\theta\in\mathbb{R}^{p}$. Training involves observations $\{(a_{i},u_{i})\}_{i=1}^{N}$ where $u_{i}=\mathcal{G}^{\dagger}(a_{i})$. The objective is to find parameters $\theta^{*}$ minimizing the loss:
@@
where $P$ maps the input function $a$ to a higher dimensional representation $f_{0}(x)$, $Q$ maps the final layer’s output $f^{(J)}(x)$ back to the target function space $\mathcal{U}$, and each intermediate layer $f^{(j+1)}$ updates the representation $f^{(j)}$ via:
$f_{l+1}(x)=\sigma\left(W^{(j)}f^{(j)}(x)+(\mathcal{K}^{(j)}(a;\psi^{(j)})f^{(j)})(x)\right),\quad\forall x\in D.$ (18)
@@
The Fourier Neural Operator (FNO) [19] proposes an efficient implementation of the kernel integral operator $\mathcal{K}^{(j)}$. The FNO restricts the kernel integration to be a convolution operator, assuming the kernel depends only on the displacement, i.e., $\kappa^{(j)}(x,y;\psi^{(j)})=\kappa^{(j)}(x-y;\psi^{(j)})$. This allows the use of the Convolution Theorem, which states that a convolution in the spatial domain is equivalent to point-wise multiplication in the Fourier domain:
@@
$f_{l+1}(x)=\sigma\left(W^{(j)}f^{(j)}(x)+\mathcal{F}^{-1}\left(R_{\psi^{(j)}}\cdot(\mathcal{F}(f^{(j)}))\right)(x)\right).$ (21)
@@ 2.3 Time-Propagator Design
To model the evolution of the electron density $n(x,t)$ under the influence of a time-dependent external potential $v_{l}(t)$, we employ an autoregressive approach based on FNOs whose architecture is well-suited for learning resolution-invariant mappings between infinite-dimensional function spaces, making it appropriate for processing spatio-temporal data like the electron density [19]. The autoregressive framework allows the model to predict future states based on a sequence of past observations and the corresponding external potential values.
@@ 2.3.1 Input Representation Design
The autoregressive FNO framework requires an effective representation of both the electron density history and the time-dependent external potential. We implement the following input representation strategies to capture the complex spatio-temporal dependencies between these quantities:
@@ 2.3.2 Observable Calculations
Total Energy: We cannot evaluate the total energy as the FNO model predicts the total electron density while we need the Kohn-Sham orbitals to evaluate the Kohn-Sham kinetic energy. Instead we use the Thomas-Fermi approximation as a proxy for the total energy functional which consists of kinetic, Hartree, and external potential contributions. The kinetic energy is approximated as:
@@
The total energy is then:
@@ 2.3.3 Physics-Informed Loss Functions
The inclusion of physics-informed losses not only improves the physical plausibility of the predictions but also enhances the model’s generalization to systems outside the training distribution.
@@ 2.3.4 Error Metrics
To evaluate model performance, we employ multiple error metrics that capture different aspects of prediction accuracy for both point-wise density fields and integrated observables.
@@
where $y_{i}$ and $\hat{y}_{i}$ denote reference and predicted values, respectively. For density fields, this is computed point-wise across all spatial locations and time steps.
@@ 3.4 Observable Predictions
Total Energy: Within the Thomas-Fermi approximation, we compute the total energy functional comprising kinetic, Hartree, and external potential contributions. Figure 3 shows the temporal evolution of predicted versus reference total energies.
@@ 4 Discussion
The FNO approach has clear limits. The demonstration is restricted to one-dimensional, two-electron diatomic molecules under ALDA and sinusoidal driving in the dipole approximation. However, our approach does generalize methodologically to three dimensions, but its feasibility needs yet to be shown. The operator advances densities directly and is not unitary at the orbital level, so constraints beyond an integral penalty are desirable. Data quality and boundary handling in the references can influence learning, even though trajectories with artefacts were filtered.

### v3

@@ 1 Introduction
Time-Dependent Density Functional Theory (TDDFT) [runge_density-functional_1984] is a widely used method to study the evolution of electronic structure under the influence of external time-dependent potentials. It is used to calculate various excited state properties such as excitation energies [adamo_calculations_2013], charge transfer [maitra_charge_2017], stopping power [yost_examining_2017], optical absorption spectra [jacquemin_excited-state_2011] and non-linear optical properties [goncharov_non-linear_2014]. Due to its favorable balance between accuracy and computational cost, TDDFT has been applied in various domains including photocatalysis [herringRecentAdvancesRealTime2023], biochemistry [varsanoTDDFTStudyExcited2006], nanoscale devices [senanayakeRealTimeTDDFTInvestigation2019] and the study of light-matter interactions in general [neufeldBenchmarkingFunctionalsStrongField2024].
@@
For weak perturbations, the linear response formalism of TDDFT is used to calculate the excitation spectrum of a system  [PGB00]. It is calculated using the Casida equation [casida_time-dependent_1995] as an eigenvalue problem that describes the first order response of the density. In contrast, the electron density is directly propagated in time under the real time formalism. Real time TDDFT can be used to calculate the nonlinear response of the density under strong perturbations such as ultrafast electron dynamics with strong laser fields [provorse_electron_2016].
@@
There are multiple components that must be decided for real time TDDFT calculations. These include preparation of the initial density and orbitals, choice of the exchange-correlation functional, the form of the external time-dependent potential which is determined by the problem to be solved and the choice of the time-propagation scheme. The time evolution is a significant fraction of the computational cost [castro_propagators_2004, gomez_pueyo_propagators_2018].
@@
Machine learning (ML) for accelerating scientific simulations is a rapidly growing area of research [carleo_machine_2019]. A variety of unsupervised and data-driven models have been developed to solve differential equations across a wide range of domains [karniadakisPhysicsinformedMachineLearning2021a].
@@
Neural operators (NOs) [kovachki_neural_2023] are a class of models that map function-to-function spaces, as opposed to the finite-dimensional vector space mappings of neural networks. This is especially useful for partial differential equation (PDE) problems where experimental or simulation data is available. Fourier neural operators (FNOs) [liFourierNeuralOperator2020] are a type of NO which represent operator weights in Fourier space. The main advantages of FNOs are that they generalize well across function spaces, and being resolution-invariant, they can be used for inference on higher resolution grids than the training set grids. FNOs have also been used for forward and inverse PDE problems in various domains [azizzadenesheli_neural_2024].
@@
While many ML models have been developed for the electronic structure in ground state DFT [fiedler_deep_2022, brockherde_bypassing_2017, eickenberg_solid_2018, schmidt_learning_2018, MiRy2019, grisafi_transferable_2019, fabrizio_electron_2019, chandrasekaran_solving_2019, tsubaki_quantum_2020, ben_mahmoud_learning_2020, lewis_learning_2021, ellis_accelerating_2021, FiMoScVo23, Rackers_2023, Shao_2023, CaFiBrSh25], there have been relatively fewer efforts focused on TDDFT. Some applications of ML for TDDFT include development of exchange-correlation potentials [ScBeMa19, suzuki_machine_2020, yang_machine-learning_2023, BhGuIs24] and predicting properties such as spectra [BoShZhXu24, ShLe25] and stopping power [WaBlLeMa24]. In addition, there is prior work on developing ML models for quantum dissipative systems [Zh25, zhangNeuralQuantumPropagators2025, zhangNeuralNetworkSolution2025] and for FNO based TDDFT time propagators for model systems under the influence of a fixed laser field [shahAcceleratingElectronDynamics2024].
@@ 2.1 Time-Dependent Density Functional Theory
where $\underline{\bm{r}}$ represents the collective coordinates of the $N$ electrons, $\Psi(\underline{\bm{r}},t)$ is the many-body wavefunction and $\hat{H}$ is the Hamiltonian operator, representing the total energy of the system. For this problem the form of $\hat{H}$ is
$\hat{H}(t)=\hat{T}+\hat{W}+\hat{V}(t),$ (2)
where $\hat{T}=\sum_{j=1}^{N}-\nabla_{j}^{2}/2$ is the kinetic energy operator, $\hat{W}=\sum_{1\leq j<k\leq N}\frac{1}{\left|\mathbf{r}_{j}-\mathbf{r}_{k}\right|}$ represents the electron-electron interaction, and $\hat{V}(t)=v_{\mathrm{{\scriptscriptstyle ext}}}(\bm{r},t)$ is the time-dependent external potential operator which includes the static external potential due to the ions and an external time-dependent potential that drives the system. Note that in writing this Schrödinger equation we are assuming the Born-Oppenheimer approximation, i.e. we are treating the dynamics of the ions as occurring on a much longer time scale than the dynamics of the electrons, and can therefore treat the ions as classical point-like particles. The initial state $\Psi(\underline{\bm{r}})=\Psi(\underline{\bm{r}},t_{0})$ is obtained by solving the time-independent Schrödinger equation
@@
which is an eigenvalue problem where $E$ denotes the eigenvalues and corresponds to the total energy of the system. The static Hamiltonian $\hat{H}$ is similar in form to Eq. (2) but with a time-independent external potential $\hat{V}$.
@@
TDDFT is based on the Runge-Gross theorem [runge_density-functional_1984] which states that for a system with a given ground state many-body wavefunction $\Psi_{0}=\Psi_{0}\left(\boldsymbol{r},t_{0}\right)$, there exists a unique mapping between the potential and the time-dependent density. The density of the interacting many-body system can be obtained by solving a system of fictitious non-interacting particles governed by the time-dependent Kohn-Sham (TDKS) equations [van_leeuwen_causality_1998]
@@
$v_{\mathrm{{\scriptscriptstyle S}}}[n](\bm{r},t)=v_{\mathrm{{\scriptscriptstyle ext}}}(\bm{r},t)+v_{\mathrm{{\scriptscriptstyle H}}}(n(\bm{r},t))+v_{\mathrm{{\scriptscriptstyle XC}}}[n](\bm{r},t),$ (7)
@@
While in theory, $v_{\mathrm{{\scriptscriptstyle XC}}}(\bm{r},t)$ depends on the density at all previous time steps, the adiabatic approximation is often used in practice [ullrich_time-dependent_2012]. We use the adiabatic local density approximation (ALDA). Under ALDA, $v_{\mathrm{{\scriptscriptstyle XC}}}(\bm{r},t)$ is approximated as:
$v_{\mathrm{{\scriptscriptstyle XC}}}^{\text{ALDA}}[n,\Psi_{0},\Phi_{0}](\mathbf{r},t)=v_{\mathrm{{\scriptscriptstyle XC}}}^{\text{LDA}}(n(\mathbf{r},t)),$ (8)
@@
These coupled equations Eq. (4) to Eq. (7) are solved through iterative numerical algorithms [castro_propagators_2004].
@@ 2.1.1 Time Propagators
Numerous methods have been developed and analyzed for approximating the short-time propagator $\hat{U}(t_{j}+\Delta t,t_{j})$ [castro_propagators_2004, gomez_pueyo_propagators_2018], including Magnus expansions that approximate the exponent in $\hat{U}(\Delta t)=\exp(\hat{\Omega}(\Delta t))$, split-operator schemes effective when the Hamiltonian separates into analytically solvable parts (e.g., kinetic and potential), explicit and implicit Runge–Kutta methods adapted to the Schrödinger equation, exponential-midpoint and other exponential integrators, and implicit schemes such as Crank–Nicolson [Crank_Nicolson_1947]. The choice of propagator often involves a trade-off between accuracy, stability (especially concerning the maximum usable time step $\Delta t$), computational cost per step, and conservation properties.
For this work, the reference data for training our neural network model was generated using the Crank-Nicolson method [Crank_Nicolson_1947]. This is an implicit method, formulated as:
@@ 2.1.2 Model Systems
For the electron-electron interaction we also consider a soft-Coulomb potential
$v_{\text{ee}}(\mathbf{r},\mathbf{r^{\prime}})=v_{\text{ee}}(x,x^{\prime})=\frac{1}{\sqrt{(x-x^{\prime})^{2}+a^{2}}}\ .$ (16)
Within the ALDA, used here for the TDDFT calculations, we adopt the one-dimensional forms of the exchange [1D_LDA_Exchange] and correlation [1D_LDA_Correlation] functionals as implemented in the Octopus real-time TDDFT code [tancogne-dejean_octopus_2020]. For both the external potential and the electron–electron interaction, the softening parameter of the soft-Coulomb interaction is chosen as $a=1$.
@@
We use the LDA functional due to its simplicity and accuracy for such simple one-dimensional systems. The density evolution is calculated by solving the time-dependent Kohn-Sham equations with the external potential $v_{\text{ext}}(x,t)$ using the Crank-Nicholson scheme. The numerical simulation is performed using the Octopus real time TDDFT code [tancogne-dejean_octopus_2020].
@@ 2.2 Physics-Informed Machine Learning
Neural operators (NOs) [luDeepONetLearningNonlinear2019, anandkumarNeuralOperatorGraph2020] extend neural networks by mapping functions to functions instead of finite-dimensional vectors. While a neural network maps $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$, a neural operator maps function spaces, $\mathcal{G}:\mathcal{A}\rightarrow\mathcal{U}$. Our goal is to approximate the non-linear map $\mathcal{G}^{\dagger}:\mathcal{A}\rightarrow\mathcal{U}$ with a neural operator $\mathcal{G}_{\theta}$, parameterized by $\theta\in\mathbb{R}^{p}$. Training involves observations $\{(a_{i},u_{i})\}_{i=1}^{N}$ where $u_{i}=\mathcal{G}^{\dagger}(a_{i})$. The objective is to find parameters $\theta^{*}$ minimizing the loss:
@@
where $P$ maps the input function $a$ to a higher dimensional representation $f_{0}(x)$, $Q$ maps the final layer’s output $f^{(J)}(x)$ back to the target function space $\mathcal{U}$. This mapping is done pointwise, projecting the vector at each spatial point into a vector of higher dimension to provide a richer latent representation for the model. Each intermediate layer $f^{(j+1)}$ updates the representation $f^{(j)}$ via:
$f_{j+1}(x)=\sigma\left(W^{(j)}f^{(j)}(x)+(\mathcal{K}^{(j)}(a;\psi^{(j)})f^{(j)})(x)\right),\quad\forall x\in D.$ (20)
@@
The Fourier Neural Operator (FNO) [liFourierNeuralOperator2020] proposes an efficient implementation of the kernel integral operator $\mathcal{K}^{(j)}$. The FNO restricts the kernel integration to be a convolution operator, assuming the kernel depends only on the displacement, i.e., $\kappa^{(j)}(x,y;\psi^{(j)})=\kappa^{(j)}(x-y;\psi^{(j)})$. This allows the use of the Convolution Theorem, which states that a convolution in the spatial domain is equivalent to point-wise multiplication in the Fourier domain:
@@
$f_{j+1}(x)=\sigma\left(W^{(j)}f^{(j)}(x)+\mathcal{F}^{-1}\left(R_{\psi^{(j)}}\cdot(\mathcal{F}(f^{(j)}))\right)(x)\right).$ (23)
@@ 2.3 Time-Propagator Design
To model the evolution of the electron density $n(x,t)$ under the influence of a time-dependent external potential $v_{l}(t)$, we employ an autoregressive approach based on FNOs whose architecture is well-suited for learning resolution-invariant mappings between infinite-dimensional function spaces, making it appropriate for processing spatio-temporal data like the electron density [liFourierNeuralOperator2020]. The autoregressive framework allows the model to predict future states based on a sequence of past observations and the corresponding external potential values.
@@ 2.3.1 Input Representation Design
The autoregressive FNO framework requires an effective representation of both the electron density history and the time-dependent external potential. The representation consists of density grids $\mathbf{N}$, external potential grids $\mathbf{V}$ and a normalized spatial grid $\mathbf{g}$ with coordinate points in the range [-1,1], providing explicit positional information. We implement the following input representation strategies to capture the complex spatio-temporal dependencies between these quantities:
@@ 2.3.2 Observable Calculations
Thomas-Fermi Total Energy: We cannot evaluate the Kohn-Sham total energy as the FNO model predicts the total electron density while we need the Kohn-Sham orbitals to evaluate the Kohn-Sham kinetic energy. Instead we use the Thomas-Fermi approximation [parr_density-functional_1995, martin_electronic_2004] as a proxy for the total energy functional which is defined as the sum of kinetic, Hartree, and external potential contributions. The kinetic energy is approximated as:
@@
The total energy in the Thomas-Fermi approximation is then:
@@ 2.3.3 Physics-Informed Loss Functions
The inclusion of physics-informed losses not only improves the physical plausibility of the predictions but also enhances the model’s generalization to systems outside the training distribution. Furthermore, we include this loss term in all models below, as inclusion of this term stabilizes training and makes models reach physical consistency faster.
@@ 2.3.4 Error Metrics
To evaluate model performance, we employ multiple error metrics that capture different aspects of prediction accuracy for both point-wise density fields and integrated observables. Here, $y_{i}$ and $\hat{y}_{i}$ denote reference and predicted values respectively.
@@
For density fields, this is computed point-wise across all spatial locations and time steps.
@@ 3.4 Observable Predictions
Thomas-Fermi Total Energy: Within the Thomas-Fermi approximation, we compute the total energy functional comprising kinetic, Hartree, and external potential contributions. Figure 3 shows the temporal evolution of predicted versus reference total energies.
@@ 4 Discussion
The FNO approach has clear limits. The demonstration is restricted to one-dimensional, two-electron diatomic molecules under ALDA and sinusoidal driving in the dipole approximation. However, our approach does generalize methodologically to three dimensions, but its feasibility needs yet to be shown. The operator advances densities directly and is not unitary at the orbital level, so constraints beyond an integral penalty are desirable. Data quality and boundary handling in the references can influence learning, even though trajectories with artefacts were filtered. While the TF approximation was used to validate energy conservation in an orbital-free manner, using orbital-dependent Kohn-Sham total energy formulation would be required for more accurate energy calculations. Extending FNO-based propagators to three-dimensional TDDFT is conceptually straightforward but computationally demanding. The FFT-based operator scales as $\mathcal{O}(n\log n)$ but the memory consumption of 3D feature tensors grows rapidly and is the primary bottleneck. Recent variants introduce architectural changes to mitigate this, enabling applications to a range of three-dimensional simulation tasks [li_geometry-informed_2023, rahman_sparsified_2024, wang_prediction_2024].